OpenStudy (anonymous):

Which logarithmic equation is equivalent to the exponential equation below? e^4x = 5 (the x is part of the ^4 exponent) A. log 5 = 4x B. ln 4x = 5 C. ln 5 = 4x D. log 4x = 5

OpenStudy (ajspeller):

Take the natural log (ln) of both sides of the equation. ln e^4x = ln 5 4x = ln 5

OpenStudy (anonymous):

I'm confused, so does the 4x=ln 5 flip and become C. or does it become B?

OpenStudy (ajspeller):

ln 5 = 4x same as .. 4x = ln 5 It's like saying 1+1=2 or that 2=1+1. Both expressions separated by the equal sign are equivalent.